Weaver Company had a net deferred tax liability of $34,000 at the beginning of the year, representing a net taxable temporary difference of $100,000 (taxed at 34%). During the year, Weaver reported pretax book income of $400,000. Included in the computation were favorable temporary differences of $50,000 and unfavorable temporary differences of $20,000. At the beginning of the year, Congress reduced the corporate tax rate to 21%. This results in a deferred tax benefit of $6700. What would be the net effect of the tax rate change to the existing balance of the deferred tax liability?
Answer #1

Favorable temporary difference = 50,000

Unfavorable temporary difference = 20,000

Net Favorable temporary difference = 50,000 - 20,000 = 30,000

Deferred income tax expense= 30,000 * 21% = 6300

Previous year's temporary difference = 100,000 * (34% - 21%) = 13000

Net deferred tax benefit due to change in tax rate = 6300 - 13000 = (6700)